Help!!!!!!!!!!!!
1.Your cousin earns about $25 per week baby-siting and recives one $5 doller bonus. You earn about $15 per week mowing lawns and $12 per week running errands. After working the same number of weeks, you have $11 more than yor cousin. write and solve an equation to find how many weeks you worked.

Answer by checkley75(3666)

25x+5+11=(15+12)x
25x+16=27x
25x-27x=-16
-2x=-16
x=-16/-2
x=8 in weeks you'll have $11 more than your cousin.
proof
25*8+16=27*8
200+16=216
216=216